-- ElenaRocco - 2011-08-04

Periods partitioning for the 2011 longitudinal run

The table will update during the data taking

Type of physics Start period run # Date and time start period End period run # Date and time end period Alignment run # Period name MD duration Intervention on detectors not during MD
Test 90948 2011-06-10 21:47 91700 2011-06-23 13:48   T23    
1T 91701 2011-06-23 13:56 91736 2011-06-24 08:06 91718/91720 W24 no MD  
2.5T 91737 2011-06-24 10:13 92168 2011-07-09 01:33 91783/91784 W25 2011-07-01/2011-07-08  
2.5T 92169 2011-07-09 01:33 92676 2011-07-27 09:08 92259/92260 and 92646 till 92650 W27 2011-07-27/2011-07-28  
2.5T 92677 2011-07-27 09:14 92869 2011-08-03 07:40 92819 till 92827 W30 2011-08-03/2011-08-04  
2.5T 92870 2011-08-03 07:45 93161 2011-08-11 07:15 93014/93015/93022/93023 (±2.5T) and 93016/93021 (±1T) W31 stop for Magnet change on SPS moved DC0 for noise investigation/vacuum gauge for Si02
2.5T 93162 2011-08-11 07:18 93325 2011-08-17 07:10 93292/93293 W32 2011-08-17/2011-08-18 during the w-e QUAD02 instability/off; no desplacement in beam position for -380A on QUAD02, but different trg rates compared to the nominal setting (-441A)
2.5T 93326 2011-08-17 07:49 93496 2011-08-20 04:40 93465/93466 (2.5T) W33 no MD 570 spills of good data. Test runs at 1 T: 93486, 93487 at +1 T, 93490 at -1 T.
2.5T 93497 2011-08-20 05:06 93735 2011-08-29 08:02 93574/ 93575 (solenoid off) W34 no MD lost polarisation; data and alignment runs w/o solenoid field taken; runs with different target thikness; test with LAS w/o CM in; alignments run taken w/o solenoid; no detectors displacement; few runs with low polarisation value
unpol. 93736 2011-08-29 09:42 94088 2011-09-09 17:05 94039/94040 W35 2011-08-29/2011-09-03 lost of polarisation:pb with solenoid
2.5T 94089 2011-09-09 18:36 94577 2011-09-21 09:46 94174/94177 W36 no MD change thr in MM, Straw; lost of polarisation during MD
2.5T 94578 2011-09-21 10:52 94844 2011-09-28 08:03 94739/94740 W38 2011-09-21/2011-09-22 ECAL2 missing;QUAD34 and TRIM07 problems;DC noise;STRAW thr changed (few channels)
2.5T 94845 2011-09-28 08:03 95287 2011-10-12 07:14 94974/94975 and 95274/95276 W39 2011-09-28 DC00 displacement (comment # 40528) lost of polarisation (comment #40652)
2.5T 95288 2011-10-12 08:27 95766 2011-10-26 07:06 95466/95467 and 95724/95726 W41 2011-12-10/2011-10-14 ECAL2 problem missing blocks
2.5T 95767 2011-10-26 08:03 96118 2011-11-7 07:43 95863/95864 and 96095/96096 W43 2011-26-10/2011-10-27 trigger test ECAL2; noise in MMs; no beam kicker pb
no magnetic field and no beam 96119 2011-11-7 08:55       T45   test after the end of the run
Edit | Attach | Watch | Print version | History: r27 < r26 < r25 < r24 < r23 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r27 - 2012-08-21 - ElenaRocco
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    Compass/SIDIS All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright &© 2008-2023 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback